OpenStudy (anonymous):

25divided by 5(2+3)=

OpenStudy (anonymous):

Your answer is 1,\[\frac{ 25 }{ 5(2+3) }\]\[\frac{ 25 }{ 5\times(5) }\]\[\frac{ 25 }{ 25 }\]\[1\]